40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jun 19 18:56:55 2018, Jon, Update, General, AUX beam alignment issues 
    Reply  Wed Jun 20 00:57:04 2018, johannes, Update, General, AUX beam alignment issues as110_rf_30MHz.pdfyend_rf_30MHz.pdf
       Reply  Sun Jun 24 23:13:46 2018, johannes, Update, General, AUX beam alignment issues AS_AUX_SETUP.pdf
Message ID: 14013     Entry time: Sun Jun 24 23:13:46 2018     In reply to: 13989
Author: johannes 
Type: Update 
Category: General 
Subject: AUX beam alignment issues 

At some point we want to change the AUX injection on the AS table to interfere less with the normal interferometer path, and avoid 10/90 beamsplitters which produce a fair amount of ghosting. The plan is to replace the 99/1 BS whose reflection goes to AS110 and AS55, while the transmission goes to the AS camera, with a 90/10 BS as shown in the attachment. This results in ~10% less light on the PDs compared to the pre-AUX era. Between this BS and the AS camera there will be a second 90/10 BS that sends the AUX light into the IFO, so we end up with marginally less AUX power into the IFO and the same PSL power on the AS cam. We're short optics, so this has to wait until two new beamsplitters arrive from CVI.

Attachment 1: AS_AUX_SETUP.pdf  2.064 MB  Uploaded Mon Jun 25 00:15:54 2018  | Hide | Hide all
AS_AUX_SETUP.pdf
ELOG V3.1.3-